OpenStudy (anonymous):

Please help!!!!! Find a value for k in the function below so that the function is continuous at all points. h(x)={(x^7+1/x+1) x≠-1 k x=-1

OpenStudy (anonymous):

Do you know what the value of k=?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

can you simplify this: \(\large \frac{x^7+1}{x+1} \) ???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do you know synthetic division?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

do you know how to do long division?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

*of polynomials??

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

ok... first i'd like you to see what the answer will be if we graph it... we can do the algebra (long division) later... here is an online graphing calculator.... can you graph: y=\(\large \frac{x^7+1}{x+1} \) ??

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

I have a graphiong calculator and I graphed it and got it.

OpenStudy (anonymous):

so what do you think k = ??? when x=-1?

OpenStudy (anonymous):

k=-1

OpenStudy (anonymous):

or k=0

OpenStudy (anonymous):

no... what is the y-value when x=-1? look at your graph....

OpenStudy (anonymous):

here's a screen shot :

OpenStudy (anonymous):

k=0

OpenStudy (anonymous):

no.... at x=-1, the y-value on the graph is not 0....

OpenStudy (anonymous):

I have an error for the y value so that is why I'm a little confused

OpenStudy (anonymous):

yes... you should have an error because the function is not defined at x=-1. but if you look at the graph, at x values close to x=-1, what does the y values seem to be close to?

OpenStudy (anonymous):

k=1

OpenStudy (anonymous):

If you want to show it without the use of a graph, you'll have to find the quotient, like @dpalnc mentioned. Then, to show continuity at x = -1, you have to find k such that \[\lim_{x\to -1}h(x)=k,\text{ i.e.}\\ \lim_{x\to -1}\frac{x^7+1}{x+1}=k\]

OpenStudy (anonymous):

so does k=-1 really confused?

OpenStudy (anonymous):

no... k is NOT -1

OpenStudy (anonymous):

ok well i'm really confused and don't get it so I NEED HELP

OpenStudy (anonymous):

Well do you know what continuity means?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

There's your problem. Brush up on the definition of continuity first, then come back.

OpenStudy (anonymous):

ok got it